Re-post: Need your help???
For the function defined by f(x) =5x-4, find a formula for f^-1(x).

Answer by venugopalramana(3286)

For the function defined by f(x) =5x-4, find a formula for f^-1(x).
LET Y=5X-4
5X=Y+4
X=(Y+4)/5
HENCE F^-1(Y)=(Y+4)/5
F^-1(X)=(X+4)/5